# Towards Safe and Reliable Autonomous Driving: Dynamic Occupancy Set Prediction

### v1

@@ IV-C Comparative Methods
  • Probability Prediction (PR)-Based Method[4, 27]: This method models the possible positions of traffic participants at each future time step as Gaussian distributions. These probabilistic models are trained using maximum likelihood estimation. To convert these probability distributions into DOSs with well-defined boundaries, the method forms ellipsoidal occupancy sets centered around the mean of the Gaussian distribution. The standard deviation of the Gaussian distribution is scaled by a factor $z$ to estimate DOSs with different ranges. Specifically, we consider three values of $z$: 0.5, 1, and 2.
@@
  • Specified Occupancy Set Method: In this method, we assume that occupancy sets are circular. The radius for each future time step is manually specified. Three variants are considered: 1) Circle-based, where a fixed radius $r$ is set for each time step; 2) Linear-based, where the radius increases linearly with time, i.e., $r=z\times t$, where $z$ is a scaling factor; 3) Error-based, where the initial radius $r_{e}$ for each time step is set based on the average error statistics, and the occupancy set radius is then $r=z\times r_{e}$.
For our proposed methods, we also implemented different variants based on the representations proposed in Section III-B: 1) DOS (circle), assuming DOS as a circle; 2) DOS (ellipse, $\theta=\theta_{0}$), assuming DOS as an ellipse with a fixed orientation after affine transformations; 3) DOS (ellipse, v1), assuming the ellipse’s orientation is determined by the current velocity direction; 4) DOS (ellipse, v2), assuming the ellipse’s orientation is determined by the predicted velocity direction at each time. 5) DOS (ellipse), where the ellipse’s orientation is estimated by the DOS prediction module.
@@ V-A Comparing Different Methods
The SA-based method, while excelling in predicting errors, may not be the most suitable for effectively balancing CR and OSA. It tends to prioritize the accuracy of predicted errors, aiming for occupancy set boundaries that closely match the average of all errors. This tendency results in nearly half of the coverage rate at $z=1$. This further highlights the necessity of our study, which focuses on achieving the specific objectives of Sufficient Coverage and Minimal DOS Area.
@@
Furthermore, we assessed the performance of our DOS methods on the SIND dataset, representing complex real-world traffic scenarios. The results, presented in Table II, reinforce our earlier observations, highlighting the adaptability and robustness of our methods across diverse scenarios. Our DOS methods consistently outperformed other approaches in the SIND dataset, demonstrating their effectiveness in challenging and dynamic traffic environments. This reaffirms their potential to ensure both safety and efficiency in autonomous driving applications.
@@ V-B Upper Limit Analysis of Different DOS Representation
In this subsection, we focus on an upper limit analysis of various DOS representations, with a primary objective of evaluating their performance under the condition of ensuring 100% coverage while minimizing DOS area. To achieve this goal, the true coordinates at each time precisely align with the DOS boundaries. For circular DOS assumptions, we determine the upper limit by setting the radius equal to the true prediction error. For elliptical DOS assumptions with specified orientation angles, we compute the upper limit by solving for the Mahalanobis distance equal to 1, thereby obtaining the theoretical minimum OSA for these representations. The analysis results are presented in Fig. 4, revealing several key observations:
@@
2) The theoretical upper limit analysis prompts a reconsideration of the difficulty and necessity of achieving complete coverage. For long-term predictions, the presence of large prediction errors at certain outlier cases may result in significantly larger required OSA, potentially compromising navigable space. Therefore, under the guidance of our optimization objective, the preference leans towards predicting DOS with relatively smaller areas while ensuring adequate coverage.
3) This theoretical analysis also provides valuable insights into future directions for this research. Firstly, by optimizing the underlying trajectory prediction algorithms, we can directly improve the theoretical upper limits. Smaller trajectory prediction errors theoretically enable achieving sufficient CR with smaller OSA. Additionally, further improvements in the practical performance of DOS (ellipse, v1), DOS (ellipse, v2), and DOS (ellipse) methods are necessary to fully exploit their potential. Furthermore, Addressing long-term DOS prediction challenges requires ongoing optimization efforts.
4) Fig. 4 includes an evaluation of the PR-based occupancy set prediction with $z=1$, providing a clear visual comparison that highlights the advantages of our proposed methods.
@@ V-C Impact of Different Parameters on DOS Results
$k$
$\alpha$
$\beta$
CR OSA
1s
4s
1s
4s
0.5
3.0
0.0
0.921
0.718
1.049
6.139
0.2
3.0
0.0
0.949
0.766
1.653
9.278
0.05
3.0
0.0
0.981
0.843
3.559
18.047
0.01
3.0
0.0
0.994
0.916
7.304
37.01
0.01
10.0
0.0
0.998
0.962
16.706
96.492
0.1
1.0
0.0
0.945
0.708
1.352
3.800
0.1
2.0
0.0
0.957
0.767
1.673
8.093
0.1
3.0
0.0
0.969
0.811
2.461
13.364
0.1
3.0
20.0
0.969
0.811
2.461
13.364
We conducted experiments to assess the impact of different parameter settings on the performance of our DOS prediction framework. Table III provides a summary of the results obtained for various combinations of $k$, $\alpha$, and $\beta$. These experiments yield valuable insights into how these parameters influence the trade-off between coverage and efficiency.
1) Impact of $k$: Our experiments revealed that smaller $k$ values, ranging from 0.5 to 0.01, led to a significant improvement in Coverage Rate (CR), with $k=0.01$ achieving a remarkable 99.4% CR at the 1-second prediction horizon. This reduction in $k$ enhances coverage; however, it tends to result in larger OSA, affecting the compactness of DOS predictions. 2)Impact of $\alpha$: Increasing the $\alpha$ values from 1.0 to 3.0 had a positive effect on CR but came at the cost of increased OSA. 3)Impact of $\beta$: The results showed that adjusting $\beta$ within the range of 0.0 to 20.0 had negligible effects on both CR and OSA. Within this tested range, $\beta$ appeared to have limited impact on the outcomes.
@@ V-D Planning with DOS
success rate collision rate out-of-bounds rate velocity
w/o DOS
0.900
0.100
0.000
5.474
w DOS
0.947
0.053
0.000
5.710

### v2

@@ IV-C Comparative Methods
  • Probability Prediction (PR)-Based Method[20]: This method models the possible positions of traffic participants at each future time step as Gaussian distributions. These probabilistic models are trained using MLE. To convert these probability distributions into DOSs with well-defined boundaries, the method forms ellipsoidal occupancy sets centered around the mean of the Gaussian distribution. The standard deviation of the Gaussian distribution is scaled by a factor $z$ to estimate DOSs with different ranges. Specifically, we consider three values of $z$: 0.5, 1, and 2.
@@
  • Specified Occupancy Set: In this method, we assume that occupancy sets are circular. The radius for each future time step is manually specified. Three variants are considered: 1) Circle-based, where a fixed radius $r$ is set for each time step; 2) Linear-based, where the radius increases linearly with time, i.e., $r=z\times t$, where $z$ is a scaling factor; 3) Error-based, where the initial radius $r_{e}$ for each time step is set based on the average error statistics, and the occupancy set radius is then $r=z\times r_{e}$.
For our proposed methods, we also implemented different variants based on the representations proposed in Section III-B: 1) DOS (circle), assuming DOS as a circle; 2) DOS (ellipse, $\theta=\theta_{0}$), assuming DOS as an ellipse with a fixed orientation after affine transformations; 3) DOS (ellipse, v1), assuming the ellipse’s orientation is determined by the current velocity direction; 4) DOS (ellipse, v2), assuming the ellipse’s orientation is determined by the predicted velocity direction at each time. 5) DOS (ellipse), where the ellipse’s orientation is also estimated by the DOS prediction module.
@@ V-A Comparing Different Methods
The SA-based method, while excelling in predicting errors, may not be the most suitable for effectively balancing CR and OSA. It tends to prioritize the accuracy of predicted errors, aiming for occupancy set boundaries that closely match the average of all errors. This tendency results in nearly half of the coverage rate at $z=1$. This further highlights the necessity of our study, which focuses on achieving the objectives of sufficient coverage and Minimal DOS Area.
@@
Furthermore, we assessed the performance of our DOS methods on the SIND dataset, representing complex real-world traffic scenarios. Due to space constraints, we presented only the most representative variant for each category of methods compared. The results, presented in Table II, reinforce our earlier observations, highlighting the adaptability and robustness of our methods across diverse scenarios. Our DOS methods consistently outperformed other approaches in the SIND dataset, demonstrating their effectiveness in challenging and dynamic traffic environments. This reaffirms their potential to ensure both safety and efficiency in autonomous driving applications.
@@ V-B Upper Limit Analysis of Different DOS Representation
In this subsection, we focus on an upper limit analysis of various DOS representations, with a primary objective of evaluating their performance under the condition of ensuring 100% coverage while minimizing DOS area. To achieve this goal, the true coordinates at each time precisely align with the DOS boundaries. For circular DOS assumptions, we determine the upper limit by setting the radius equal to the true prediction error. For elliptical DOS with specified orientation angles, we compute the upper limit by constraining the Mahalanobis distance to equal 1, thereby deriving the theoretical minimum OSA for these representations. The results are shown in Fig. 4, which elucidates several insights:
@@
2) The upper limit analysis prompts a reconsideration of the difficulty and necessity of achieving complete coverage. For long-term predictions, the presence of large prediction errors at certain outlier cases may result in significantly larger required OSA, potentially compromising navigable space. Therefore, under the guidance of our optimization objective, the preference leans towards predicting DOS with relatively smaller areas while ensuring adequate coverage.
3) This analysis also provides valuable insights into future directions for this research. Firstly, by optimizing the underlying trajectory prediction algorithms, we can directly improve the theoretical upper limits. Smaller trajectory prediction errors theoretically enable achieving sufficient CR with smaller OSA. Additionally, further improvements in the practical performance of DOS (ellipse, v1), DOS (ellipse, v2), and DOS (ellipse) methods are necessary to fully exploit their potential. Furthermore, Addressing long-term DOS prediction challenges requires ongoing optimization efforts.
4) Fig. 4 includes an evaluation of the PR-based method with $z=1$, providing a clear visual comparison that highlights the advantages of our proposed methods.
@@ V-C Impact of Different Parameters on DOS Results
$k$ $\alpha$ CR (%) OSA
1s 4s 1s 4s
0.5 3.0 92.1 71.8 1.049 6.139
0.2 3.0 94.9 76.6 1.653 9.278
0.05 3.0 98.1 84.3 3.559 18.047
0.01 3.0 99.4 91.6 7.304 37.01
0.01 10.0 99.8 96.2 16.706 96.492
0.1 1.0 94.5 70.8 1.352 3.800
0.1 2.0 95.7 76.7 1.673 8.093
0.1 3.0 96.9 81.1 2.461 13.364
We conducted experiments to assess the impact of different parameter settings on the performance of our DOS prediction framework. Table III provides a summary of the results obtained for various combinations of the key parameters $k$ and $\alpha$. These experiments yield valuable insights into how these parameters influence the trade-off between coverage and efficiency: 1) Impact of $k$: Our experiments revealed that smaller $k$ values, ranging from 0.5 to 0.01, led to a significant improvement in Coverage Rate (CR), with $k=0.01$ achieving a remarkable 99.4% CR at the 1-second prediction horizon. This reduction in $k$ enhances coverage; however, it tends to result in larger OSA, affecting the compactness of DOS predictions. 2)Impact of $\alpha$: Increasing the $\alpha$ values from 1.0 to 3.0 had a positive effect on CR but came at the cost of increased OSA.
@@ V-D Planning with DOS
success rate (%) collision rate (%) out-of-bounds rate (%) velocity
w/o DOS 90.0 10.0 0.0 5.474
w DOS 94.7 5.3 0.0 5.710